The function H described by H(x)= 2.75x + 71.48 can be used to estimate the height, in centimeters, of a woman whose humerus is x cm long. Estimate the height of a woman whose humerus is 34.4 cm long.
Found 2 solutions by uma, rapaljer:
Answer by uma(370)

H(x)= 2.75x + 71.48
Given the humerus of a woman = 34.4 cm
So her ht = 2.75(34.4) + 71.48
= 166.08
Thus the ht of the woman = 166.08 cm

Answer by rapaljer(4671)

Find H(x)= 2.75x + 71.48 , where x=34.4 cm.
There is that "Glade" commercial again: "Plug it in, plug it in!!" You take the formula that is given, and substitute (in English we say "plug it in") the value of x, and find the value of H.
H(x) = 2.75x+ 71.48, so
H(34.4) = 2.75*34.4 + 71.48 (hopefully you have a calculator??)
H(34.4) = 166.08 cm.

R^2 at SCC